A brand name has an 80% recognition rate. If the owner of the brand wants to verify that rate by beginning with a small sample of 5 randomly selected consumers, find the probability that exactly 4 of the 5 consumers recognize the brand name. Also find the probability that the number who recognize the brand is not 4.
Answer by stanbon(75887)

A brand name has an 80% recognition rate. If the owner of the brand wants to verify that rate by beginning with a small sample of 5 randomly selected consumers,
----------------------
Binomial Problems with n = 5 and p(recognize) = 0.8
------
Find the probability that exactly 4 of the 5 consumers recognize the brand name.
P(x = 4) = 5C4(0.8)^4*(0.2) = binompdf(5,0.8,4) = 0.4096
---------------------------
Also find the probability that the number who recognize the brand is not 4.
P(x # 4) = 1 - 0.4096 = 0.5604
====================
Cheers,
Stan H.
====================